OpenStudy (anonymous):

Show an equation and a solution for the problem. Crane A can unload a ship in 10 hours and crane B can unload it in 14 hours. How long will it take the two cranes to unload the ship working together? Round to the nearest tenth.

OpenStudy (anonymous):

\[\frac{10\times 14}{10+14}\] is the number you are looking for

OpenStudy (anonymous):

crane A has rate \[\frac{1}{10}\] per hour and crane B has rate \[\frac{1}{14}\] their combined rate is therefore \[\frac{1}{10}+\frac{1}{14}=\frac{10+14}{10\times 14}\] and you want to solve \[\frac{10+14}{10\times 14}T = 1\] for T so the answer is \[\frac{10\times 14}{10+14}\]